# Simple Policy Optimization

### v4

@@ Abstract
PPO (Proximal Policy Optimization) algorithm has demonstrated excellent performance in many fields, and it is considered as a simple version of TRPO (Trust Region Policy Optimization) algorithm. However, the ratio clipping operation in PPO may not always effectively enforce the trust region constraints, this can be a potential factor affecting the stability of the algorithm. In this paper, we propose Simple Policy Optimization (SPO) algorithm, which introduces a novel clipping method for KL divergence between the old and current policies. Extensive experimental results in Atari 2600 environments indicate that, compared to the mainstream variants of PPO, SPO achieves better sample efficiency, extremely low KL divergence, and higher policy entropy, and is robust to the increase in network depth or complexity. More importantly, SPO maintains the simplicity of an unconstrained first-order algorithm. Code is available at https://github.com/MyRepositories-hub/Simple-Policy-Optimization.
Machine Learning, ICML
@@ 1 Introduction
Policy based model-free reinforcement learning algorithms have been widely applied in various fields, and have achieved performance surpassing humans including board games (Silver et al., 2016, 2017, 2018), video games (Berner et al., 2019; Vinyals et al., 2019; Ye et al., 2020), autonomous driving (Kiran et al., 2021; Teng et al., 2023) and intelligent control (Chatzilygeroudis et al., 2019; Yu & Wang, 2022; Wen et al., 2022). Although there are two main learning paradigms in deep reinforcement learning: value-based and policy-based, impressive algorithms such as PPO (Schulman et al., 2017) and A3C (Mnih et al., 2016) rely on the Actor-Critic (AC) framework (Konda & Tsitsiklis, 1999), which is a synthesis of the two paradigms.
Figure 1: We over-optimize the objective in the Amidar-v5 environment for 100 epochs and select five random seeds for experiments. SPO-$x$ represents the SPO algorithm with the maximum KL divergence set to $x$, PPO-$x$ represents the PPO algorithm with the ratio clipping parameter set to $x$, and SO represents directly optimizing the surrogate objective.
@@
Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm implicitly limits the difference between the old policy and the current policy through the ratio clipping operation, which is reflected in the Trust Region Policy Optimization (TRPO) (Schulman et al., 2015a) algorithm as a hard constraint on their KL divergence expectation. Ratio clipping avoids solving the expensive KL divergence constraint problem, so the PPO algorithm can also be well extended to large-scale reinforcement learning environments (Berner et al., 2019; Ye et al., 2020). Prior works have discussed how ratio clipping can effectively implement trust region constraints. For example, even if the probability ratio is bounded, the corresponding KL divergence is not necessarily bounded (Wang et al., 2020). The performance improvement of PPO algorithm is likely to mainly depend on “code-level optimizations” (Engstrom et al., 2020). There are also some works that design novel probability ratio clipping methods or objective functions (Queeney et al., 2021; Garg et al., 2021; Markowitz & Staley, 2023; Chen et al., 2018), or alternate between a standard RL phase and a distillation phase during the training stage to improve sample efficiency of PPO (Cobbe et al., 2021; Wang et al., 2023). There are many ways to avoid excessive KL divergence, which gives birth to many variants of PPO. For example, KL divergence can be added as a penalty term to the surrogate objective and introduce adaptive penalty coefficient (Schulman et al., 2017). During the training process of PPO, a KL threshold is set, and an early stopping strategy is adopted when the KL divergence reaches the threshold (Achiam, 2018). Alternatively, an early stopping strategy can be implemented based on the deviation degree of probability ratio instead of KL divergence (Sun et al., 2022). However, the above works may not have fundamentally changed the characteristics of the PPO algorithm (we acknowledge that the surrogate objective with ratio clipping remains the preferred choice in most cases), nor have they effectively constrained the KL divergence between the old and current policies. We propose Simple Policy Optimization (SPO) algorithm to compensate for these gaps, SPO does not even require the clipping of probability ratio, instead of a novel clipping method for the KL divergence between the old and current policies. Using SPO, we highlight three important observations about it:
@@ 2.1 Policy Gradient
Reinforcement learning is generally defined by a tuple $(\mathcal{S},\mathcal{A},r,\mathcal{P},\rho_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space, $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the probability distribution of the state transition function, $\rho_{0}:\mathcal{S}\mapsto\mathbb{R}$ is the initial state distribution, and $\gamma\in[0,1]$ is the discount factor. For each trajectory
@@
$J(\theta)=\mathbb{E}_{\tau\sim p(\cdot)}\left[R(\tau)\right]=\sum_{\tau}R(\tau)\cdot p(\tau),$ (2)
where $R(\tau)=\sum_{i=1}^{T-1}\gamma^{i-1}r_{i}$, the policy gradient can be estimated as
$\nabla J(\theta)\approx\frac{1}{N}\sum_{n=1}^{N}\sum_{t=1}^{T_{n}-1}R(\tau^{n})\cdot\nabla\log\pi_{\theta}(a_{t}^{n}|s_{t}^{n}).$ (3)
$\nabla J(\theta)$ can be estimated through the Monte Carlo (MC) method (Williams, 1992).
@@ 2.2 Trust Region Policy Optimization
Theorem 2.1.
(Schulman et al., 2015a) Let $\alpha=D_{\rm TV}^{\max}(\pi,\tilde{\pi})$ and $\epsilon=\max_{s}\lvert\mathbb{E}_{a\sim\tilde{\pi}}\left[A^{\pi}(s,a)\right]\rvert$, the following bound holds:
$\eta(\tilde{\pi})\geq L_{\pi}(\tilde{\pi})-\frac{2\epsilon\gamma}{(1-\gamma)^{2}}\cdot\alpha^{2},$
where $L_{\pi}$ matches $\eta$ to first order if the policy $\pi$ is a differentiable function of a parameter vector (Kakade & Langford, 2002).
Noting the relationship between total variation divergence and the KL divergence $D_{\rm TV}(\pi,\tilde{\pi})^{2}\leq D_{\rm KL}(\pi,\tilde{\pi})$, we have the following corollary.
Corollary 2.2.
@@
$\eta(\tilde{\pi})\geq L_{\pi}(\tilde{\pi})-\frac{2\epsilon\gamma}{(1-\gamma)^{2}}\cdot D_{\rm KL}^{\max}(\pi,\tilde{\pi}).$
Through a heuristic approximation, the maximum KL divergence $D_{\rm KL}^{\max}$ in the Corollary 2.2 is approximated as the average KL divergence and used as a constraint to obtain the formalization of TRPO algorithm, that is,
@@
1:  Input: Policy network parameters $\theta$; value network parameters $w$; threshold of KL divergence $d_{\max}$
@@
8:        Use GAE (Schulman et al., 2015b) technique to estimate the advantage $\hat{A}(s_{t},a_{t})$
@@
10:        For each data $(s_{t},a_{t},r_{t})$ calculate $d\leftarrow\sum_{a\in\mathcal{A}}\pi_{\theta_{\rm old}}(a|s_{t})\cdot\log\frac{\pi_{\theta_{\rm old}}(a|s_{t})}{\pi_{\theta}(a|s_{t})}$ and $d_{\text{clip}}\leftarrow\text{clip}(d,0,d_{\max})$
@@
$l_{p}\leftarrow\left\{\begin{aligned} &-\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t}),&\enspace\theta=\theta_{\rm old}\\ &-\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t})\cdot\left\{\frac{d_{\rm clip}}{d}+\text{sign}\left[\hat{A}(s_{t},a_{t})\right]-1\right\}\cdot\text{sign}\left[\hat{A}(s_{t},a_{t})\right],&\enspace\theta\neq\theta_{\rm old}\\ \end{aligned}\right.$
@@
2.3 Proximal Policy Optimization
$\begin{split}\max_{\theta}\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\Big{\{}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}&\cdot\hat{A}(s_{t},a_{t})\\ -\beta&\left[D_{\text{KL}}(\pi_{\theta_{\text{old}}},\pi_{\theta})\right]\Big{\}},\\ \end{split}$ (5)
$\begin{split}\max_{\theta}\enspace&\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t}),\\ &\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon\right)\cdot\hat{A}(s_{t},a_{t})\Bigg{]}\Bigg{\}}.\\ \end{split}$ (6)
3 Method
In TRPO, trust region are imposed by limiting the expectation of the KL divergence between the old and current policies. Here, we denote the KL divergence as
$\begin{split}d&=D_{\text{KL}}\left[\pi_{\theta_{\rm old}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right]\\ &=\sum_{a\in\mathcal{A}}\pi_{\theta_{\rm old}}(a|s_{t})\cdot\log\frac{\pi_{\theta_{\rm old}}(a|s_{t})}{\pi_{\theta}(a|s_{t})},\\ \end{split}$ (7)
and the surrogate objective
According to the Corollary 2.2, we hope to optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ with the KL divergence $d$ as small as possible, as it optimizes the monotonic improvement lower bound in policy iterations. This minimax problem seems to be equivalent to the following form:
[…]
where
$\sigma=\text{sign}\left[\hat{A}(s_{t},a_{t})\right]=\left\{\begin{aligned} +1&,&\enspace\hat{A}(s_{t},a_{t})>0;\\ 0&,&\enspace\hat{A}(s_{t},a_{t})=0;\\ -1&,&\enspace\hat{A}(s_{t},a_{t})<0.\\ \end{aligned}\right.$ (13)
[…]
(a) Training curves of return.
(b) Training curves of average KL divergence.
(c) Training curves of policy entropy.

### v5

@@ Abstract
PPO (Proximal Policy Optimization) algorithm has demonstrated excellent performance in many fields, and it is considered as a simple version of TRPO (Trust Region Policy Optimization) algorithm. However, the ratio clipping operation in PPO may not always effectively enforce the trust region constraints, this can be a potential factor affecting the stability of the algorithm. In this paper, we propose Simple Policy Optimization (SPO) algorithm, which introduces a novel clipping method for KL divergence between the old and current policies. Extensive experimental results in Atari 2600 environments indicate that, compared to the mainstream variants of PPO, SPO achieves better sample efficiency, extremely low KL divergence, and higher policy entropy, and is robust to the increase in network depth or complexity. More importantly, SPO maintains the simplicity of an unconstrained first-order algorithm. Our code is available at https://github.com/MyRepositories-hub/Simple-Policy-Optimization.
@@ 1 Introduction
Policy based model-free reinforcement learning algorithms have been widely applied in various fields, and have achieved performance surpassing humans including board games [1, 2, 3], video games [4, 5, 6], autonomous driving [7, 8] and intelligent control [9, 10, 11]. Although there are two main learning paradigms in deep reinforcement learning: value-based and policy-based, impressive algorithms such as PPO [12] and A3C [13] rely on the Actor-Critic (AC) framework [14], which is a synthesis of the two paradigms.
@@
Proximal Policy Optimization (PPO) [12] algorithm implicitly limits the difference between the old policy and the current policy through the ratio clipping operation, which is reflected in the Trust Region Policy Optimization (TRPO) [15] algorithm as a hard constraint on their KL divergence expectation. Ratio clipping avoids solving the expensive KL divergence constraint problem, so the PPO algorithm can also be well extended to large-scale reinforcement learning environments [4, 6]. Prior works have discussed how ratio clipping can effectively implement trust region constraints. For example, even if the probability ratio is bounded, the corresponding KL divergence is not necessarily bounded [16]. The performance improvement of PPO algorithm is likely to mainly depend on “code-level optimizations” [17]. There are also some works that design novel probability ratio clipping methods or objective functions [18, 19, 20, 21], or alternate between a standard RL phase and a distillation phase during the training stage to improve sample efficiency of PPO [22, 23].
There are many ways to avoid excessive KL divergence, which gives birth to many variants of PPO. For example, KL divergence can be added as a penalty term to the surrogate objective and introduce adaptive penalty coefficient [12]. During the training process of PPO, a KL threshold is set, and an early stopping strategy is adopted when the KL divergence reaches the threshold [24]. Alternatively, an early stopping strategy can be implemented based on the deviation degree of probability ratio instead of KL divergence [25].
However, the above works may not have fundamentally changed the characteristics of the PPO algorithm (we acknowledge that the surrogate objective with ratio clipping remains the preferred choice in most cases), nor have they effectively constrained the KL divergence between the old and current policies. We propose Simple Policy Optimization (SPO) algorithm to compensate for these gaps, SPO does not even require the clipping of probability ratio, instead of a novel clipping method for the KL divergence between the old and current policies. Using SPO, we highlight three important observations about it:
@@ 2.1 Policy Gradient
Reinforcement learning is generally defined by a tuple $(\mathcal{S},\mathcal{A},r,\mathcal{P},\rho_{0},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space, $r:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the probability distribution of the state transition function, $\rho_{0}:\mathcal{S}\mapsto[0,1]$ is the initial state distribution, and $\gamma\in[0,1]$ is the discount factor.
For each trajectory
@@
$J(\theta)=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[R(\tau)\right]=\sum_{\tau}R(\tau)\cdot p_{\theta}(\tau),$ (2)
where $\theta$ is the parameters of the policy network, $R(\tau)=\sum_{i=1}^{T-1}\gamma^{i-1}r_{i}$, and the policy gradient can be estimated as
$\begin{split}\nabla J(\theta)&=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[R(\tau)\cdot\nabla\log p_{\theta}(\tau)\right]\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\log\left[\rho_{0}(s_{1})\cdot\prod_{t=1}^{T-1}\pi_{\theta}(a_{t}|s_{t})\cdot\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left\{R(\tau)\cdot\nabla\left[\log\rho_{0}(s_{1})+\sum_{t=1}^{T-1}\log\pi_{\theta}(a_{t}|s_{t})+\log\mathcal{P}(s_{t+1}|s_{t},a_{t})\right]\right\}\\ &=\mathbb{E}_{\tau\sim p_{\theta}(\cdot)}\left[\sum_{t=1}^{T-1}R(\tau)\cdot\nabla\log\pi_{\theta}(a_{t}|s_{t})\right]\approx\frac{1}{N}\sum_{n=1}^{N}\sum_{t=1}^{T_{n}-1}\left[R(\tau^{n})\cdot\nabla\log\pi_{\theta}(a_{t}^{n}|s_{t}^{n})\right],\\ \end{split}$ (3)
now we can estimate $\nabla J(\theta)$ through the Monte Carlo (MC) method [26].
@@ 2.2 Trust Region Policy Optimization
Theorem 1.
[15] Let $\alpha=D_{\rm TV}^{\max}(\pi,\tilde{\pi})$ and $\epsilon=\max_{s}\lvert\mathbb{E}_{a\sim\tilde{\pi}}\left[A^{\pi}(s,a)\right]\rvert$, the following bound holds:
$\eta(\tilde{\pi})\geq L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot\alpha^{2},$ (4)
where $L_{\pi}$ matches the original objective $\eta$ to first order if the policy $\pi$ is a differentiable function of a parameter vector [27].
Noting the relationship between total variation divergence and the KL divergence $D_{\rm TV}(\pi,\tilde{\pi})^{2}\leq D_{\rm KL}(\pi,\tilde{\pi})$, we have the following corollary:
Corollary 2.
@@
$\eta(\tilde{\pi})\geq L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot D_{\rm KL}^{\max}(\pi,\tilde{\pi}).$ (5)
Theoretically, this lower bound is really powerful, this is because if we denote it as
$M_{\pi}(\tilde{\pi})=L_{\pi}(\tilde{\pi})-\frac{4\epsilon\gamma}{(1-\gamma)^{2}}\cdot D_{\rm KL}^{\max}(\pi,\tilde{\pi}),$ (6)
we have $\eta(\tilde{\pi})\geq M_{\pi}(\tilde{\pi})$ and $\eta(\pi)=M_{\pi}(\pi)$, and then we obtain
$\eta(\tilde{\pi})-\eta(\pi)\geq M_{\pi}(\tilde{\pi})-M_{\pi}(\pi).$ (7)
This means that we can indirectly optimize the original objective $\eta$ by optimizing the lower bound $M_{\pi}$. Through a heuristic approximation, the maximum KL divergence $D_{\rm KL}^{\max}$ in the Corollary 2 is approximated as the average KL divergence and used as a constraint to obtain the formalization of TRPO algorithm, that is,
@@
2.3 Proximal Policy Optimization
$\max_{\theta}\enspace\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\left\{\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t})-\beta\left[D_{\text{KL}}(\pi_{\theta_{\text{old}}},\pi_{\theta})\right]\right\},$ (9)
$\max_{\theta}\enspace\mathop{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta_{\rm old}}}\Bigg{\{}\min\Bigg{[}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})}\cdot\hat{A}(s_{t},a_{t}),\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\rm old}}(a_{t}|s_{t})},1\pm\epsilon\right)\cdot\hat{A}(s_{t},a_{t})\Bigg{]}\Bigg{\}}.$ (10)
3 Method
In TRPO, trust region are imposed by limiting the expectation of the KL divergence between the old and current policies. Here, we denote the KL divergence under state $s_{t}$ as
$d=D_{\text{KL}}\left[\pi_{\theta_{\rm old}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right]=\sum_{a\in\mathcal{A}}\pi_{\theta_{\rm old}}(a|s_{t})\cdot\log\frac{\pi_{\theta_{\rm old}}(a|s_{t})}{\pi_{\theta}(a|s_{t})},$ (11)
and the surrogate objective
According to the Corollary 2, we hope to optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ with the KL divergence $d$ as small as possible, as it optimizes the monotonic improvement lower bound in policy iterations. This minimax problem seems to be equivalent to the following form:
However, note that if we simply using the above form as the objective for the policy network, this can lead to significant problems. Now Assume that the estimated value of the advantage $\hat{A}(s_{t},a_{t})$ is positive, the policy network would strive to minimize the KL divergence in order to maximize the objective. This tendency would result in the network preferring to make only minor adjustments to the parameters, thereby hindering its ability to learn effectively from historical data.
[…]
We will provide intuition about the objective (16), suppose that $d_{\max}=0.01$, and the function curves is shown in Figure 1. If $\hat{A}(s_{t},a_{t})>0$ and $d>d_{\max}$, the objective $J(\theta)$ will decrease rapidly (see curve $d_{\text{clip}}/d$), this can force the KL divergence to satisfy the trust region constraints. Otherwise, when $\hat{A}(s_{t},a_{t})<0$ and $d>d_{\max}$, the function curve $-d_{\text{clip}}/d+2$ and $d_{\text{clip}}/d$ are symmetric about the line $y=1$, and $J(\theta)$ will also decrease. Therefore, to maximize the objective $J(\theta)$, SPO will secretly optimize the surrogate objective $\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})$ within the trust region $d\leq d_{\max}$ as much as possible, but it is still an unconstrained first-order algorithm.
[…]
Considering that the hyperparameter $d_{\max}$ can be set manually and even dynamically changed at any moment during the training process, SPO can benefit significantly from the Corollary 2. In summary, the characteristics mentioned above enable SPO to potentially achieve better monotonic policy improvement than PPO, and its performance depends on the value of the hyperparameter $d_{\max}$, which corresponds to the degree of conservativeness in policy iterations.
@@ 3 Method
1:  Input: Policy and value network parameters $\theta,w$; threshold of KL divergence $d_{\max}$
@@
8:        Use GAE [28] technique to estimate the advantage $\hat{A}(s_{t},a_{t})$
@@
10:        For each $(s_{t},a_{t})$ calculate $d\leftarrow D_{\text{KL}}\left[\pi_{\theta_{\rm old}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right]$ and $d_{\text{clip}}\leftarrow\text{clip}(d,0,d_{\max})$
@@
$l_{p}\leftarrow\left\{\begin{aligned} &-\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t}),\enspace&\theta=\theta_{\rm old}\\ &-\mathfrak{O}_{\theta_{\rm old}}^{\theta}(s_{t},a_{t})\cdot\left(\frac{d_{\rm clip}}{d}+\sigma-1\right)\cdot\sigma,\enspace&\theta\neq\theta_{\rm old}\\ \end{aligned}\right.$
@@
4 Experiments
We report results on the Atari 2600 environment [29, 30]. The Atari 2600 environment is a popular benchmark in the field of reinforcement learning. It consists of a set of video game environments designed for the Atari 2600 gaming console. These environments serve as testing grounds for developing and evaluating reinforcement learning algorithms and have been widely used because of its diverse set of games, ranging from simple classics like Pong to more complex titles like SpaceInvaders and Breakout. This diversity allows researchers to assess the generalization capabilities of reinforcement learning algorithms across different tasks and complexities.
We preprocessed the Atari 2600 environment, for example, the input to the network is the stacking of the last four frames resized to $84\times 84\times 4$, reward clipping and so on [31, 32]. In all experiments, we use the hyperparameters provided in the Appendix A unless otherwise specified. For each environment in Atari 2600, we compute and visualize the standard deviation of the algorithm’s performance across 3 separate runs with different random seeds. Our experiments are divided into five parts:
  1. 1.
    Comparison to PPOs. The first part compares the performance of several PPO variants and the SPO algorithm in the Atari 2600 environment, as well as the changes in KL divergence.
  2. 2.
    Over optimization. In the second part, we increase the number of update epochs from 8 to 1000 in a certain training stage to over-optimize the respective objectives of PPO and SPO, and observe their optimization of surrogate objectives and the growth trend of KL divergence.
  3. 3.
    Increasing network complexity. In the third part, we try to increase the depth of the neural network and observe how the complexity of the network structure affects the performance of PPO and SPO in the training process.
  4. 4.
    The scalability of SPO for continuous action space. In the fourth part, we further tested the scalability of SPO in a continuous action space environment, e.g., MuJoCo [33].
  5. 5.
    Sensitivity analysis. Finally, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of the SPO algorithm. Due to the limitation of space, please refer to the Appendix B.7 for the detailed content of this section.